Write the equation of the line that satisfies the given conditions. Express the final equation in standard form.
Contains the point (-4, 9) and is parallel to the line x − 2y = 6

Answer by Alan3354(69443)

Write the equation of the line that satisfies the given conditions. Express the final equation in standard form.
Contains the point (-4, 9) and is parallel to the line x - 2y = 6
--------------
Find the slope of the line line x - 2y = 6
-----
x - 2y = 6
y = (1/2)x - something, doesn't matter.
Slope m = 1/2
All parallel lines have the same slope.
---------
Use y = mx + b and the point (-4,9) to find b, the y-intercept.
9 = (1/2)*-4 + b = b-2
b = 11
-----------
y = (1/2)x + 11
x - 2y = -11 standard form
Notice only the coefficient is different.